An insulated cup contains 330 g of water at 20 °C. A 100 g sample of an unknown material at 76 °C is added. The final temperature of the water and metal is 47 °C. 

Specific heat capacity of water = 4160 J/kg-°C


What is the specific heat capacity of the unknown material {_1} J/kg-°C


1
Expert's answer
2022-05-23T10:18:02-0400

"c_2=c_1\frac{m_1}{m_2}\frac{t-t_1}{t_2-t}=12800~\frac{J}{kg\cdot °C}."